# Revisiting Privacy, Utility, and Efficiency Trade-offs when Fine-Tuning Large Language Models

### v1

@@ 5 Comparison of Fine-tuning Methods
In this section, we compare all fine-tuning methods across three key aspects: privacy, utility, and efficiency. Recall that privacy is measured as training loss on sensitive tokens and utility is measured as test loss on non-sensitive tokens. For efficiency, we measure floating point operations (FLOPs) based on the number of operations incurred (e.g., matrix multiplication, addition, etc.) during training. In this section, for DP we use a noise ratio of $\sigma=0.1$ and for LoRA we use $r=16,\alpha=16$. Figures 14(a)-14(c) present the Pareto plots for CustomerSim, comparing the three fine-tuning strategies across three dimensions: privacy, utility, and efficiency.
Privacy: Regarding privacy, as previously noted, full fine-tuning shows poor privacy performance over extended training, while DP achieves the highest privacy levels. LoRA provides a similar degree of privacy compared to DP throughout most training epochs for small-scale models (Pythia and Gemma), but declines with the large-scale model (Llama2).
@@
Utility: In terms of utility, full fine-tuning maintains relatively strong performance throughout most of its training (right on the x-axis). DP while yielding good utility in smaller models (Pythia and Gemma), performs poorly in the larger model (Llama2) and overall needs more epochs to reach higher utility levels. In contrast, LoRA consistently preserves higher utility across the entire training period.
@@
Efficiency: The color bar in Figure 14 highlights the FLOPs intensity associated with each model across all fine-tuning strategies. DP requires the highest number of FLOPs (in red) due to the need for per-sample gradient computation, where each sample corresponds to a token within each training sequence. FFT demands a moderate number of FLOPs (in blue), proportional to the total number of parameters. Finally, LoRA requires the fewest FLOPs (in green), as especially during backpropagation, most operations only operate on the low-rank matrices.
@@
3. LoRA, a parameter-efficient fine-tuning method, maintains high utility and achieves privacy levels comparable to DP in smaller models, though this advantage reduces in larger ones. Figure 14(c) shows that while LoRA preserves less privacy as training progresses, it is possible to select checkpoints that balance strong privacy with utility. This finding challenges the prevailing notion that privacy must come at the cost of high efficiency, demonstrating that LoRA can offer privacy benefits. Moreover, LoRA-tuned models retain performance on benchmark datasets close to that of pre-trained models throughout training. Figure 5 in Appendix 16 shows a comprehensive comparison across all the fine-tuning methods.
@@ 7 Ethics considerations
This study utilized publicly available datasets [44, 18], some of which included identifiable information such as personal details. However, third-party organizations pre-processed and validated the datasets to ensure that no real individuals’ data were present, thus mitigating potential privacy concerns.
@@ 11 Prompt for Annotating Privacy-Sensitive Information
We used the following prompt to obtain the privacy-sensitive information from GPT-4o for CustomerSim: {mdframed}[backgroundcolor=lavender, linewidth=0pt] “You are a private data identifying bot. You will be provided a conversation between a user and a customer service bot. You need to identify ALL instances of private data in the conversation. Private data includes any information that can be used to identify an individual, such as names, phone numbers, email addresses, and locations. It can also include any other sensitive information, such as credit card numbers, social security numbers, tracking data, health information etc.. Identify all of these use your discretion to identify any other information that can be used to identify an individual.”
12 Human Survey: Assessing Privacy-Sensitivity in Memorized Segments
We conducted a human annotation survey to assess if instances memorized by an LLM contain privacy-sensitive information, sampling 100 sequences divided into 5 questionnaires. Each was given to 40 Prolific workers. After an introductory explanation, participants evaluated each instance by (1) flagging privacy-sensitive information (ignoring routine commands unless uniquely identifiable) and (2) listing detected items if flagged. Figure 6 shows that most questions were labeled as non-sensitive.
@@ 13 Human Survey: Assessing the Quality of Privacy-Sensitive Annotations from Presidio and GPT-4
We evaluated privacy-sensitive annotations from Presidio and GPT-4 through two Prolific surveys, each with 10 samples for 40 participants. Binary Survey: For each sample, participants viewed side-by-side annotations from both tools (order randomized) and chose the more accurate one. Figure 17(a) shows GPT-4 was consistently preferred over Presidio for identifying privacy-sensitive information. Multiple-Choice Survey: Each of the 10 instances included 5 samples from each tool, displayed randomly. Participants had four options: (1) Accurate (all sensitive info correctly annotated), (2) Under-Annotated (some sensitive info missed), (3) Over-Annotated (non-sensitive info included), and (4) Mixed-Annotated (both missed and non-sensitive info annotated). Figure 17(b) shows GPT-4 rated as “Accurate” across datasets.
@@
14 Why distinguish between sensitive and non-sensitive entities?
Figure 18 shows the plots observed in Figure 2 in log-scale for better distinction.
(a) Privacy measure in Pythia
(b) Utility measure in Pythia
(c) Privacy measure in Gemma
(d) Utility measure in Gemma
(e) Privacy measure in Llama2
(f) Utility measure in Llama2
@@ 15 Exploring the Tradeoffs between Privacy and Utility
We see a similar trade-off between utility and privacy when using Presidio annotations in Figure  19, as observed in Figure 7. For the CustomerSim dataset, we notice that privacy decreases at the beginning of training, while utility improves. However, as training progresses, both metrics start to worsen. In contrast, the SynBio dataset shows a decline in both utility and privacy, except for the Pythia model, which initially experiences an improvement in utility during the first few epochs.
@@
A similar pattern as in Figure 9 using GPT-4 annotations, has also emerged in figure 20 using Presidio annotations. We observe that in the CustomerSim dataset 20, DP maintains privacy effectively with minimal degradation. However, when it comes to utility, we notice that lower noise values lead to better utility, but also result in a decrease in privacy. The same holds true for the SynBio dataset 21, with the only exception being Gemma experiencing a decline in utility after a few iterations.
@@
Comparing the results from Figures 22 and 23 using Presidio annotations with Figure 11, which used GPT-4 annotations, we can observe the same trend. This suggests that the findings are consistent across different annotation methods. Similarly, when we analyze the trade-off for the SynBio dataset in Figures  24 and 25, we see similar observations to those discussed in Figure 12, using GPT-4 annotations.
@@
Analyzing the trade-offs between privacy, utility, and efficiency in Figures 26 and 27 for CustomerSim and SynBio using Presidio annotations reveals the same efficiency ranking observed in Figure 14 with GPT-4 annotations. LoRA is the most efficient method, while DP is the least efficient, requireing the highest number of FLOPs.

### v2

@@ 5 Comparison of Fine-tuning Methods
In this section, we compare all fine-tuning methods along three key dimensions—privacy, utility, and efficiency—using the best hyperparameter configurations identified in Section 4. We present Pareto-optimal curves under a fair setting, where each method is evaluated at its best configuration. By examining these curves, one can select the checkpoint most appropriate for the task at hand.
Recall that utility is measured as test loss on non-sensitive tokens. Besides measuring privacy as the training loss on sensitive tokens, we define two more metrics for privacy : (a) Privacy Loss over sensitive tokens [10.1145/2976749.2978318] and (b) Canary exposure [carlini2019secret].
Privacy Loss: Let $\mathcal{D}$ be the dataset with the sensitive token $d$ under consideration, and $\mathcal{D^{\prime}}$ be the dataset without it. Considering $\mathcal{M}(\mathcal{D})$ as the fine-tuned model that has seen the datapoint $d$, $\mathcal{M}(\mathcal{D^{\prime}})$ as the model that has not seen $d$, and leveraging the definition from [10.1145/2976749.2978318], we can define the privacy loss (PL) as :
$\begin{split}PL&=log\frac{P(\mathcal{M}(\mathcal{D})=d)}{P(\mathcal{M}(\mathcal{D^{\prime}})=d)}=log\frac{P_{\mathcal{D}}(d)}{P_{\mathcal{D^{\prime}}}(d)}\end{split}$ (4)
A natural approximation of such a model with the data unseen is the pretrained model. Recall that negative log-likelihood for a token $d$ is $NLL_{\mathcal{(}D)}(d)=-logP_{\mathcal{D}}(d)$. We may rewrite the privacy loss as the difference between the negative log-likelihood of the pretrained model and the fine-tuned model over the token $d$ :
$\begin{split}PL&=log\frac{P_{\mathcal{D}}(d)}{P_{\mathcal{D^{\prime}}}(d)}=logP_{\mathcal{D}}(d)-logP_{\mathcal{D^{\prime}}}(d)\\ &=NLL_{\mathcal{D^{\prime}}}(d)-NLL_{\mathcal{D}}(d)\end{split}$ (5)
Unlike our measure where we observe the absolute loss on sensitive tokens from the training data, this metric observes the relative change in loss w.r.t a model that has never seen the datapoint.
Canary exposure: Originally proposed by [carlini2019secret] to measure unintended memorization, canaries are random sequences (e.g., ss=“My ID is $\circ\circ\circ\circ\circ$") inserted into training data. Exposure is computed by enumerating all possible sequences from the randomness space $\mathcal{R}$ and calculating the negative log-rank. Following the definition of exposure in [carlini2019secret], the exposure of a canary $s[r]$ in a model $\mathcal{M}$ over randomness space $\mathcal{R}$ is defined as:
$exposure_{\mathcal{M}}=log_{2}|\mathcal{R}|-log_{2}rank_{\mathcal{M}}(s[r])$ (6)
where the rank of a canary is its index in the list of all possibly-instantiated canaries, ordered by the model perplexity of all those sequences. In this setting, we inserted the canary “My ID is 34175" into the training data for $10$ times and measured its exposure as a metric for privacy.
For efficiency, we measure floating point operations (FLOPs) based on the number of operations incurred (e.g., matrix multiplication, addition, etc.) during training.
(a) Pythia
(b) Gemma
(c) Llama2
(d) Qwen2.5
Figure 14: Pareto-optimal tradeoffs (top-row : Training loss on sensitive tokens ; middle row : Privacy Loss on sensitive tokens ; bottom-row : Canary exposure). DP-SGD achieves best privacy-utility tradeoffs on small models with reduced utility on large models, albeit with a significantly high computational cost; FFT offers reasonable efficiency-utility tradeoffs with the worst privacy; and LoRA provides a balanced trade-off among all objectives.
We obtain the best configuration for DP-SGD with noise ratio of $\sigma=0.1$, and for LoRA with $r=\alpha=16,\sigma=16$. Figure 14 presents the pareto-optimal curves for CustomerSim, comparing the fine-tuning strategies over all the metrics.
@@
Privacy: Regarding privacy, FFT shows poor privacy over extended training over all 3 metrics, while DP achieves the highest privacy levels. LoRA provides similar privacy as DP throughout most epochs but declines gradually with extended training, especially on the large-scale model. Utility: FFT maintains relatively strong utility on CustomerSim. DP-SGD while yielding good utility in smaller models (Pythia,Gemma), performs poorly in the larger models (Llama2,Qwen2.5). In contrast, LoRA consistently preserves higher utility across the entire training period.
@@
Efficiency: The color bar in Figure 14 highlights the FLOPs intensity associated with each fine-tuning strategy. DP-SGD requires the highest number of FLOPs (in red) due to the need for per-sample gradient computation, where each sample corresponds to a token within each training sequence. FFT demands a moderate number of FLOPs (in blue), proportional to the total number of parameters. Finally, LoRA requires the fewest FLOPs (in green), as especially during backpropagation, most operations only operate on the low-rank matrices.
@@
3. LoRA, a parameter-efficient fine-tuning method, maintains high utility and achieves privacy levels comparable to DP in smaller models, though this advantage reduces in larger ones. Figure 14 shows that while LoRA preserves less privacy as training progresses, it is possible to select checkpoints that balance strong privacy with utility. This finding challenges the prevailing notion that privacy must come at the cost of high efficiency, demonstrating that LoRA can offer privacy benefits. Moreover, LoRA-tuned models retain performance on benchmark datasets close to that of pre-trained models throughout training. Figure 5 in Appendix 17 shows a comprehensive comparison across all the fine-tuning methods.
6 Conjecture: LoRA’s Privacy Benefits Relative to DP-SGD
In this section, we present a formal conjecture on LoRA’s privacy benefits, drawing an analogy to differential privacy (DP-SGD).
Figure 16: Privacy improves with low-rank matrix.
Analogy between LoRA and DP-SGD: We hypothesize that LoRA’s low-rank constraint helps in preserving privacy of the sensitive data. By forcing all the parameter updates into a low-rank subspace, LoRA can reduce the influence a training point can have on the model. Essentially, this is the same quantity that DP-SGD also aims to reduce by clipping and adding noise to the gradients. DP-SGD follows the route of adding randomness, and LoRA follows the route of compressing updates. The effect becomes more on the sensitive data as they are rarely occurring in the training set, which may possibly lead to it being absent in the subspace learnt by LoRA. Empirically, Figure 16 shows that low-rank LoRA model results in better privacy than its higher rank counterparts or full fine-tuning. We formally show the analogy in the theorem below:
Theorem 1.
Let $D$ and $D^{\prime}$ be two neighbouring datasets differing by a datapoint $i$. Both LoRA and DP-SGD reduce the influence of $i$ in comparison to a fully trained model, i.e. $\Delta_{DP}^{i}\leq\Delta_{FFT}^{i}$ and $\Delta_{LoRA}^{i}\leq\Delta_{FFT}^{i}$. This indicates privacy preservation of datapoint $i$ as less the influence, less will be its distinguishability.
Proof.
Consider we have $D$ and $D^{\prime}$ as two neighbouring datasets differing by a datapoint $i$. Let $W(D)$ and $W(D^{\prime})$ be the respective model parameters where $W(.)\in\mathbb{R}^{d\times k}$. Similarly, let $\Delta W(D)$ and $\Delta W(D^{\prime})$ be the change in respective aggregate model gradients. Since $D$ and $D^{\prime}$ differ by a point $i$, we can attribute the difference between $\Delta W(D)$ and $\Delta W(D^{\prime})$ to the contribution or influence of datapoint $i$ on the aggregate update.
[…]
For DP-SGD, it is well established that noise (of scale $\sigma$) gets added to the clipped gradients (gradients clipped by a threshold of $T$) that masks the impact of the datapoint [10.1145/2976749.2978318]. We can thereby frame the influence of the datapoint $i$ w.r.t DP-SGD model as:
[…]
$\therefore\Delta_{DP}^{i}\leq\Delta_{FFT}^{i}$
When LoRA with low rank $r$ is applied, the weight updates get projected to a subspace $\mathcal{S}$ of rank $r$. Let the projection be $P_{r}:\mathbb{R}^{d\times k}\rightarrow\mathcal{S}_{r}$. We can frame the influence of the datapoint $i$ w.r.t LoRA model as:
[…]
$\therefore\Delta_{LoRA}^{i}\leq\Delta_{FFT}^{i}$
We can see that both LoRA and DP-SGD reduce the influence of the datapoint ; while DP-SGD does it probabilistically using noise $\sigma$, LoRA does it deterministically through low-rank projection $\mathcal{P}_{r}$. This aligns with prior work [chaudhuri2013near, kapralov2013differentially] linking principal component analysis (PCA) or low-rank factorization to DP, where PCA limits leakage by discarding components (similar as in LoRA) and DP-SGD by adding noise.
@@ 8 Ethics considerations
This study utilized publicly available datasets [shi-etal-2022-selective, pii], some of which included identifiable information such as personal details. However, third-party organizations pre-processed and validated the datasets to ensure that no real individuals’ data were present, thus mitigating potential privacy concerns.
@@ 9 Open science
Overview of Appendices
  • Appendix 10: Limitations
  • Appendix 11: Use of LLMs
  • Appendix 12: Hyperparameters
  • Appendix 13: Validating GPT-4 predictions
  • Appendix 14: Prompt for annotating data.
  • Appendix 15: Assessing quality of annotations through human survey
  • Appendix 16: Privacy Utility Efficiency tradeoffs
  • Appendix 17: Overall comparison
10 Limitations
One of the limitations of our work lies in fine-tuning the models for unsupervised setup and not extending it to other supervised tasks like question-answering and so on. However we believe that our findings would hold in any task setup, given that the nature of these fine-tuning techniques would not change. We experimented with LoRA as one the most generic PEFT methods – however, testing out other PEFT methods would be an interesting extension of our work to explore privacy benefits extensively in the systems community.
We consider the sensitive data to be from PIIs defined in GDPR Article 4(1) [EU2016GDPR]. However, sensitive data can also be context-dependent. While one piece of information may be sensitive for one, it may not be for the other,. However, we also acknowledge that while contextual integrity is an important task, it is still hard to formalize or implement as seen in [shvartzshnaider2025position]. Lastly, while this work provides an empirical measure of privacy across different fine-tuning methods, one can definitely use such a measure for optimisation during training and establish a theoretical bound on the privacy benefits that would then also be empirically validated. We intend to explore these directions in the future.
11 Use of LLMs for Paper-writing
In this paper, we use LLMs for the following purposes:
  1. 1.
    Polishing the writing: We check for grammatical mistakes, and make minor para-phrasing to improve the flow and coherence of the paper.
  2. 2.
    Related Work: Besides traditional search, we use the OpenAI Deep Research to identify relevant literature.
@@ 14 Prompt for Annotating Privacy-Sensitive Information
We used the following prompt to obtain the privacy-sensitive information from GPT-4o for CustomerSim: “You are a private data identifying bot. You will be provided a conversation between a user and a customer service bot. You need to identify ALL instances of private data in the conversation. Private data includes any information that can be used to identify an individual, such as names, phone numbers, email addresses, and locations. It can also include any other sensitive information, such as credit card numbers, social security numbers, tracking data, health information etc.. Identify all of these use your discretion to identify any other information that can be used to identify an individual.”
@@ 15 Human Survey: Assessing the Quality of Privacy-Sensitive Annotations from Presidio and GPT-4
We evaluated privacy-sensitive annotations from Presidio and GPT-4 through two Prolific surveys, each with 10 samples for 40 participants. Figure 18 shows sample screenshots of the conducted survey. Binary Survey: For each sample, participants viewed side-by-side annotations from both tools (order randomized) and chose the more accurate one. Figure 19(a) shows GPT-4 was consistently preferred over Presidio for identifying privacy-sensitive information. Multiple-Choice Survey: Each of the 10 instances included 5 samples from each tool, displayed randomly. Participants had four options: (1) Accurate (all sensitive info correctly annotated), (2) Under-Annotated (some sensitive info missed), (3) Over-Annotated (non-sensitive info included), and (4) Mixed-Annotated (both missed and non-sensitive info annotated). Each was given to 40 Prolific workers. Figure 19(b) shows GPT-4 rated as “Accurate” across datasets.
(a) Instructions for annotators
(b) Guidelines for annotators
(c) Sample Question 1
(d) Sample Question 2
Figure 18: Annotators were provided with these instructions and guidelines to mark the samples as accurately annotated, under-annotated, over-annoted or mixed-annotated.
@@ 16 Exploring the Tradeoffs between Privacy and Utility
Full Fine-tuning: We see a similar trade-off between utility and privacy when using Presidio annotations in Figure  20, as observed in Figure 7. For the CustomerSim dataset, we notice that privacy decreases at the beginning of training, while utility improves. However, as training progresses, both metrics start to worsen. In contrast, the SynBio dataset shows a decline in both utility and privacy, except for the Pythia model, which initially experiences an improvement in utility during the first few epochs.
DP-SGD : A similar pattern as in Figure 9 using GPT-4 annotations, has also emerged in Figure 21 using Presidio annotations. We observe that in the CustomerSim dataset, DP maintains privacy effectively with minimal degradation. However, when it comes to utility, we notice that lower noise values lead to better utility, but also result in a decrease in privacy. The same holds true for the SynBio dataset in Figure 22, with the only exception being Gemma experiencing a decline in utility after a few iterations.
LoRA : Comparing the results from Figures 23 and 24 using Presidio annotations with Figure 11, which used GPT-4 annotations, we can observe the same trend. This suggests that the findings are consistent across different annotation methods. Similarly, when we analyze the trade-off for the SynBio dataset in Figures  25 and 26, we see similar observations as we observed in Figure 12 using GPT-4 annotations.
We analyze the trade-offs between privacy, utility, and efficiency for different privacy measures over SynBio with both GPT4 and Presidio annotations. Figures 27 , 28 for CustomerSim and SynBio using Presidio annotations and our privacy measure reveals the same as observed in Figure 14 with GPT-4 annotations. We also show the same for the other privacy measures – (a) privacy loss in Figures 29, 30, and 31 , and (b) exposure in Figure 32.
The overall takeaway holds consistent – LoRA is the most efficient method, followed by FFT while DP is the least efficient, requiring the highest number of FLOPs. For privacy, DP and LoRA perform at par with FFT leaking the most amount of information. In terms of utility, LoRA and FFT perform better while DP deteriorates on large scale models. Additionally, LoRA also maintains benchmark performance across other models as shown in Figure 33.
@@
(a) Pythia
(b) Gemma
(c) Llama2
(d) Qwen2.5
Figure 29: Full fine-tuning, DP and LoRA on CustomerSim with Presidio annotations for privacy loss
(a) Pythia
(b) Gemma
(c) Llama2
(d) Qwen2.5
Figure 30: Full fine-tuning, DP and LoRA on SynBio with GPT4 annotations for privacy loss
(a) Pythia
(b) Gemma
(c) Llama2
(d) Qwen2.5
Figure 31: Full fine-tuning, DP and LoRA on SynBio with Presidio annotations for privacy loss
(a) Pythia
(b) Gemma
(c) Llama2
(d) Qwen2.5
Figure 32: Full fine-tuning, DP and LoRA on CustomerSim with Presidio annotations for canary exposure.
(a) Pythia : SCIQ
(b) Pythia : HellaSwag
(c) Gemma : SCIQ
(d) Gemma : HellaSwag
(e) Llama2 : SCIQ
(f) Llama 2 : HellaSwag
Figure 33: LoRA demonstrates knowledge retention on benchmark datasets throughout the training regime on CustomerSim for other models as well, while FFT and DP show fragility with sharp and gradual performance declines, respectively.